# Vevo2: Bridging Controllable Speech and Singing Voice Generation via Unified Prosody Learning

### v1

@@ Abstract
Controllable human voice generation, particularly for expressive domains like singing, remains a significant challenge. This paper introduces Vevo2, a unified framework for controllable speech and singing voice generation. To tackle issues like the scarcity of annotated singing data and to enable flexible controllability, Vevo2 introduces two audio tokenizers: (1) a music-notation-free prosody tokenizer that captures prosody and melody from speech, singing, and even instrumental sounds, and (2) a low-frame-rate (12.5 Hz) content-style tokenizer that encodes linguistic content, prosody, and style for both speech and singing, while enabling timbre disentanglement. Vevo2 consists of an auto-regressive (AR) content-style modeling stage, which aims to enable controllability over text, prosody, and style, as well as a flow-matching acoustic modeling stage that allows for timbre control. Particularly, during pre-training of the AR model, we propose both explicit and implicit prosody learning strategies to bridge speech and singing voice. Moreover, to further enhance the AR model’s ability to follow text and prosody, we design a multi-objective post-training task that integrates both intelligibility and prosody similarity alignment. Experimental results show that the unified modeling in Vevo2 brings mutual benefits to both speech and singing voice generation. Additionally, Vevo2’s effectiveness across a wide range of synthesis, conversion, and editing tasks for both speech and singing further demonstrates its strong generalization ability and versatility. Audio samples are are available at https://versasinger.github.io/. We will release code and model checkpoints at Amphion [^1] https://github.com/open-mmlab/Amphion Zhang et al. (2024a).
1 Introduction
@@
Controllable generation of human voice is an important research topic in the field of audio generation Tan (2023); Xie et al. (2024). In recent years, remarkable progress has been achieved in speech generation, particularly in zero-shot text-to-speech (TTS) Ju et al. (2024); Anastassiou et al. (2024a); Wang et al. (2025a); Du et al. (2024a), largely propelled by the availability of large-scale training corpora. However, for more challenging domains such as singing voice—which can be regarded as a highly expressive form of speech with strictly constrained prosody (e.g., melodySundberg and Rossing (1990)—achieving both high generation quality and controllability remains challenging Bai et al. (2024); Yuan et al. (2025).
@@
This paper focuses on unifying controllable speech and singing voice generation. Our motivation is based on the hypothesis that learning for speech and singing voice can be mutually beneficial through unified modeling. Specifically, the abundance of speech data can supplement the typically smaller singing voice datasets He et al. (2024); Zhang et al. (2024b), potentially improving the overall quality of generated singing voice Yuan et al. (2025). Conversely, the inherent expressiveness of singing voice, when incorporated into a unified learning framework, can enhance expressive speech generation Gu et al. (2024) and strengthen prosody-following capabilities, which is especially advantageous for tasks such as speech editing (Section 4.3).
However, building such a unified system poses significant challenges. First, corpora for singing voice generation have traditionally relied heavily on expert annotations Umbert et al. (2015); Liu et al. (2022); Zhang et al. (2024c); Zhao et al. (2025), such as detailed music notation with precise alignment between lyrics and musical notes Zhang et al. (2022a); Dai et al. (2023); Zhang et al. (2024b). These datasets are scarce, which severely limits scalability. Moreover, this unique data organization is not well-suited for unified modeling of speech and singing voice. Second, how can we achieve precise controllability within this unified system, enabling independent manipulation of text (or lyrics), prosody (or melody), style (such as accent, emotion, or singing techniques Wilkins et al. (2018); Zhang et al. (2024b)), and timbre (i.e., speaker identity)?
To address these challenges, this study first proposes two distinct audio tokenizers: (1) Prosody Tokenizer: A single-codebook VQ-VAE tokenizer van den Oord et al. (2017); Kumar et al. (2023) operating at 6.25 Hz (56.25 bps), trained to reconstruct the chromagram Müller (2015) of raw audio. This tokenizer unifies the modeling of both melody in singing voices and prosody in speech, and can directly extract tokens from the chromagram of raw audio, eliminating the need for expert-annotated music notation and thus greatly improving scalability. Besides, due to the chromagram’s strong generalization ability across various types of audio data Müller (2015); Copet et al. (2023), this tokenizer can even capture the melody of non-human sounds, such as instrumental music (Section 4.3), thereby enhancing the flexibility of melody control in downstream controllable music generation tasks Bai et al. (2024); Yuan et al. (2025). (2) Content-Style Tokenizer: It employs a similar architecture to the prosody tokenizer, but operates at 12.5 Hz (175 bps) and is trained to reconstruct both chromagram features and Whisper’s hidden features Radford et al. (2023). It effectively captures linguistic content, melody, and style in both speech and singing voice, while also achieving robust timbre disentanglement (Table 3). Its role is similar to the “semantic tokens” in speech generation Borsos et al. (2023); Wang et al. (2025a); Du et al. (2024a); Zhang et al. (2025a), but it applies to both speech and singing voice, and operates at a lower frame rate of 12.5 Hz.
Building upon these proposed tokenizers, we introduce Vevo2, a versatile and controllable voice generation framework designed to unify speech and singing voice. Vevo2 adopts a two-stage architecture, similar to classic zero-shot speech generation systems Anastassiou et al. (2024a); Du et al. (2024a); Zhang et al. (2025a) (Figure 1). It includes an Auto-Regressive (AR) content-style modeling stage, which can be prompted by a style reference, and a Flow-Matching (FM) acoustic modeling stage, which can be prompted by a timbre reference. The uniqueness of Vevo2 lies in its AR stage: (1) Pre-training: To enable unified prosody learning for speech and singing voice, we design two different training strategies: Explicit Prosody Learning (EPL), where prosody tokens are provided as explicit input, allowing the model to predict content-style tokens from both text and prosody tokens; and Implicit Prosody Learning (IPL), where prosody tokens are not used as input, and content-style tokens are predicted solely from text tokens. By randomly applying both EPL and IPL to all voice data, the model learns prosody in a more unified manner, effectively bridging the gap between speech and singing characteristics. (2) Post-training: To further improve the text- and prosody-following abilities of the pre-trained model, we introduce a multi-objective post-training task. We propose a composite reward that incorporates both intelligibility and prosody similarity, and extend the GRPO algorithm Shao et al. (2024) for its optimization. Post-training significantly enhances the model’s controllability with respect to text and melody, and also improves its generalization to out-of-pretraining-distribution data, such as instrumental sounds.
The contributions of this paper are summarized as follows:
  • We propose two audio tokenizers: (1) a unified prosody tokenizer capable of encoding prosody and melody from speech, singing voice, and even instruments sounds, without reliance on music notation; and (2) a low-frame-rate (12.5 Hz) content-style tokenizer that models linguistic content, prosody, and style for both speech and singing voice while achieving timbre disentanglement.
  • We introduce Vevo2, a unified and controllable framework for both speech and singing voice generation. Its effectiveness across diverse synthesis, conversion, and editing tasks demonstrates strong generalization and versatility. Moreover, its superior melody-following capability enables unique music generation applications, such as humming-to-singing and instrument-to-singing.
  • We propose a multi-objective post-training task for controllable human voice generation. By optimizing both intelligibility and prosody similarity rewards with a GRPO-based algorithm, Vevo2 achieves greater stability and improved overall performance across a variety of tasks.
2 Related Work
Controllable Voice Generation Controllable speech and singing voice generation has consistently remained a significant research focus. (1) In the speech domain, existing works explore control over multiple attributes: linguistic content, which is fundamental to TTS Tan (2023); Xie et al. (2024) and speech editing tasks Le et al. (2023); Peng et al. (2024); Wang et al. (2025b); Chen et al. (2025); timbre, which is central to VC tasks Sun et al. (2016); Qian et al. (2019); Ju et al. (2024); Anastassiou et al. (2024b); Zhang et al. (2025a); Liu (2024); and style, which encompasses accent Zhao et al. (2021); Xue et al. (2024); Zhang et al. (2025a), emotion Zhou et al. (2022); Qi et al. (2024); Zhang et al. (2025a), whisper Meenakshi and Ghosh (2018), and a broader natural language description Lyth and King (2024); Ji et al. (2024); Diwan et al. (2025). (2) For singing voice, beyond controlling lyrics, timbre, and style—as explored in SVS Liu et al. (2022); Zhang et al. (2022b, 2024c, 2024d); Guo et al. (2025), SVC Huang et al. (2023); Ning et al. (2023); Sha et al. (2024); Zhang et al. (2024e), and singing lyric editing tasks Zhang et al. (2022c); Hong et al. (2023); Yang et al. (2025)melody control holds particular significance due to the inherent musicality of singing voices Copet et al. (2023); Bai et al. (2024); Yuan et al. (2025). (3) From a unified modeling perspective, previous research has attempted to integrate various speech generation tasks such as TTS and VC Zhang et al. (2025a); Du et al. (2024a); Wang et al. (2025c), or different singing voice generation tasks like SVS and SVC Hong et al. (2023); Dai et al. (2025). However, models capable of unifying both speech and singing voice domains remain scarce Yang et al. (2024a); Lei et al. (2023). Moreover, the potential benefits of joint speech and singing pre-training, along with the versatile control mechanisms for multiple voice attributes within a unified system, remain largely unexplored.
Speech Tokenizer Recent research has emphasized the disentangled speech representation learning, aiming to separately model the factorized speech attributes to reduce learning complexity Jiang et al. (2023); Ju et al. (2024) and achieve more precise controllability Polyak et al. (2021); Ju et al. (2024); Zhang et al. (2025a). Common speech tokenizers include linguistic content tokenizers Hsu et al. (2021); Zhang et al. (2025a); Baade et al. (2025), content-style tokenizers Borsos et al. (2023); Anastassiou et al. (2024a); Wang et al. (2025a); Du et al. (2024a); Zhang et al. (2025a), and prosody tokenizers Jiang et al. (2023); Ju et al. (2024). However, in pursuing a unified controllable model for both speech and singing generation, existing speech tokenizers exhibit several limitations: As these tokenizers predominantly focus on the speech domain, their effectiveness in handling singing voice data remains unexplored. The design of a unified prosody tokenizer capable of handling both modalities, along with enhancing the melody encoding capabilities of content-style tokenizers Zhang et al. (2024e), remains an important issue for future exploration.
Post-training for Audio Generation Alignment via post-training has demonstrated its effectiveness in the generation of text Ouyang et al. (2022); Bai et al. (2022), vision Xu et al. (2023); Fu et al. (2024), speech Zhang et al. (2024f); Anastassiou et al. (2024a); Du et al. (2024b), music Cideron et al. (2024), and sound effects Majumder et al. (2024); Liao et al. (2024). In speech generation, existing works have employed preference alignment to enhance multiple aspects of speech, including intelligibility Anastassiou et al. (2024a); Du et al. (2024b); Zhang et al. (2025b), speaker similarity Anastassiou et al. (2024a); Du et al. (2024b), emotion controllability Anastassiou et al. (2024a); Gao et al. (2024), and overall quality Zhang et al. (2024f); Chen et al. (2024); Yao et al. (2025). For developing a unified speech and singing voice generation model, the importance of specific capability elicitation through post-training becomes particularly crucial, given the expanded range of capabilities injected into the pre-trained model. However, the design of appropriate alignment objectives for unified speech and singing voice generation remains an open research problem.
3 Methodology
@@
    We denote the text tokens, prosody tokens, content-style tokens, and acoustic representations (specifically, Mel spectrogram) as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$, and $\boldsymbol{\bm{mel}}$, respectively. We use blue, red, and green to indicate the source inputs of these representations. $\boldsymbol{\bm{\hat{cs}}}$ and $\boldsymbol{\bm{\hat{mel}}}$ represent the model-predicted outputs. For the singing voice synthesis (SVS) task, we render MIDI notations into instrumental sounds and then extract their prosody tokens (Section 4.3).
@@
3.1 Prosody Tokenizer
Motivation In order to model both the prosody of speech and the melody of singing voice using a unified method, we choose chromagram feature Müller (2015) and design a VQ-VAE tokenizer van den Oord et al. (2017); Kumar et al. (2023) based on it. Some previous studies of speech generation tried to use fundamental frequency (F0) Polyak et al. (2021); Ju et al. (2024) or the low-frequency band of the mel spectrogram Jiang et al. (2023) to build a prosody tokenizer. The reasons why we adopt the chromagram are as follows: (1) Octave-free, thus friendly for unification: There is a significant distribution gap of F0 between speech and singing voice, such as the F0 range of singing voice is much wider Zhang et al. (2023). Chromagram removes octave information between singing voice and speech, making it easier to capture common prosodic patterns across both domains. (2) Notation-free, thus friendly for scaling up: Chromagram effectively represents high-level musical and singing voice attributes such as melody and harmony Müller (2015), and can be automatically extracted from audio without manual transcription or music notation, making it suitable for large-scale data processing. (3) Robust for diverse audio domains: As a statistical distribution of spectral energy, chromagram is less sensitive to noise and pitch detection errors De Cheveigné and Kawahara (2002), resulting in more stable performance. In addition, chromagram provides strong representations for both human voice and non-human sounds, such as instrumental music Müller (2015); Copet et al. (2023). These characteristics enhance the flexibility of melody control in downstream music generation tasks, such as enabling instrument-to-sing applications (Section 4.4).

### v2

# Vevo2: A Unified and Controllable Framework for Speech and Singing Voice Generation

@@ Abstract
Controllable human voice generation, particularly for expressive domains like singing, remains a significant challenge. This paper introduces Vevo2, a unified framework for controllable speech and singing voice generation. To tackle issues like the scarcity of annotated singing data and to enable flexible controllability, Vevo2 introduces two audio tokenizers: (1) a unified music-notation-free prosody tokenizer that captures prosody and melody from speech, singing, and even instrumental sounds, and (2) a unified content-style tokenizer that encodes linguistic content, prosody, and style for both speech and singing, while enabling timbre disentanglement. Vevo2 consists of an auto-regressive content-style modeling stage, which aims to enable controllability over text, prosody, and style, as well as a flow-matching acoustic modeling stage that allows for timbre control. Particularly, during the speech-singing joint training of the AR model, we propose both explicit and implicit prosody learning strategies to bridge speech and singing voice. Moreover, to further enhance the Vevo2’s ability to follow text and prosody, we design a multi-objective post-training task that integrates both intelligibility and prosody similarity alignment. Experimental results show that the unified modeling in Vevo2 brings mutual benefits to both speech and singing voice generation. Additionally, Vevo2’s effectiveness across a wide range of synthesis, conversion, and editing tasks for both speech and singing further demonstrates its strong generalization ability and versatility. Audio samples are are available at https://versasinger.github.io/.
I Introduction
@@
Controllable generation of human voice is an important research topic in the field of audio generation [tts-book-tanxu, xietianxin-survey]. In recent years, remarkable progress has been achieved in speech generation, particularly in zero-shot text-to-speech (TTS) [naturalspeech3, seedtts, maskgct, cosyvoice], largely propelled by the availability of large-scale training corpora. However, for more challenging domains such as singing voice—which can be regarded as a highly expressive form of speech with strictly constrained prosody (e.g., melody) [^1] This paper adopts a definition in which prosody refers to the suprasegmental attributes of human voice, including pitch, duration, and loudness. We regard melody—which primarily involves pitch and duration (with rhythm viewed as a temporal organization of duration)—as a specific and more constrained form of prosody. [singingvoice-science]—achieving both high generation quality and controllability remains challenging [seedmusic, yue].
@@
This paper focuses on unifying controllable speech and singing voice generation. Our motivation is based on the hypothesis that learning for speech and singing voice can be mutually beneficial through unified modeling. Specifically, (1) Speech $\rightarrow$ Singing Voice: the abundance of speech data can supplement the typically smaller singing voice datasets [emilia, gtsinger], potentially improving the overall quality of generated singing voice [yue]. (2) Singing Voice $\rightarrow$ Speech: the inherent expressiveness of singing voice, when incorporated into a unified learning framework, can enhance expressive speech generation [yicheng-trans]. Besides, since singing voices inherently contain structured melodic patterns, learning to model them helps the system enhance its ability to follow melody, thereby improving overall prosody-following capability. This capability is particularly beneficial for tasks such as prosody-preserved speech editing (Section V-B).
Several recent studies have attempted to unify speech and singing voice generation. For instance, UniSyn [unisyn] integrates TTS and Singing Voice Synthesis (SVS) through a task identifier that switches between generation modes, while UniAudio [uniaudio] extends this concept to a broader range of audio generation tasks via task-conditioned sequence modeling. Although these approaches represent valuable progress, they still face two fundamental challenges. First, the existing corpora for singing voice generation rely heavily on expert annotations [singing-expression, diffsinger, tcsinger, spsinger], such as detailed music notation with precise alignment between lyrics and musical notes [m4singer, singstyle111, gtsinger]. These datasets are scarce, and their unique data organization is inherently inconsistent with large-scale speech corpora, thereby limiting scalability and unified training. Second, achieving versatile controllability within such a unified system remains nontrivial—such as the independent manipulation of text (or lyrics), prosody (or melody), style (e.g., accent, emotion, or singing techniques [vocalset, gtsinger]), and timbre (i.e., speaker identity) within a single framework.
@@
    We denote the text tokens, prosody tokens, content-style tokens, and acoustic representations (specifically, Mel spectrogram) as $\boldsymbol{\bm{t}}$, $\boldsymbol{\bm{p}}$, $\boldsymbol{\bm{cs}}$, and $\boldsymbol{\bm{mel}}$, respectively. We use blue, red, and green to indicate the source inputs of these representations. $\boldsymbol{\bm{\hat{cs}}}$ and $\boldsymbol{\bm{\hat{mel}}}$ represent the model-predicted outputs. For the singing voice synthesis (SVS) task, we render MIDI notations into instrumental sounds and then extract their prosody tokens (Section V-B).
@@
To address these challenges, this paper introduces Vevo2, a versatile and controllable voice generation framework designed to unify speech and singing voice. Vevo2 adopts a two-stage architecture, similar to classic zero-shot speech generation systems [seedtts, cosyvoice, vevo] (Figure 1): (1) Auto-Regressive (AR) Content-Style Modeling Stage: given a target text and a prosodic source as input, we utilize an AR transformer to generate the content-style tokens, which is prompted by a style reference. (2) Flow-Matching (FM) Acoustic Modeling Stage: given content-style tokens as input, we employ an FM transformer to produce acoustic representations (specifically, Mel spectrograms in our study), which is prompted by a timbre reference. To achieve the desired unification and controllability, Vevo2 introduces the following three key method designs:
1. Unified Speech and Singing Voice Tokenizers. We propose two distinct audio tokenizers: a unified prosody tokenizer and a unified content-style tokenizer. (1) Prosody Tokenizer: A single-codebook VQ-VAE tokenizer [vq-vae, rvqgan] operating at 6.25 Hz (56.25 bps), trained to reconstruct the chromagram [meinard-book] of raw audio (Figure 2(a)). This tokenizer unifies the modeling of both melody in singing voices and prosody in speech, and can directly extract tokens from the chromagram of raw audio, eliminating the need for expert-annotated music notation and thus greatly improving scalability. Moreover, due to the chromagram’s strong generalization ability across various types of audio data [meinard-book, musicgen], this tokenizer can even capture the melody of non-human sounds such as instrumental music (Section V-F), thereby enabling unique music generation applications, such as humming-to-singing and instrument-to-singing. (2) Content-Style Tokenizer: It employs a similar architecture to the prosody tokenizer, but operates at 12.5 Hz (175 bps) and is trained to reconstruct both chromagram features and Whisper hidden features [whisper] (Figure 2(b)). It effectively captures linguistic content, melody, and style in both speech and singing voice, while also achieving robust timbre disentanglement (Table III).
2. Speech-Singing Joint Training. To address the joint training of speech and singing voice, we design two different training strategies during the content-style modeling stage: (1) Explicit Prosody Learning (EPL), where prosody tokens are provided as explicit input, allowing the model to predict content-style tokens from both text and prosody tokens; and (2) Implicit Prosody Learning (IPL), where prosody tokens are not used as input, and content-style tokens are predicted solely from text tokens. By randomly applying both EPL and IPL to all voice data, the model learns prosody in a more unified manner, effectively bridging the gap between speech and singing characteristics.
3. Multi-objective Alignment. To further improve the text- and prosody-following abilities of the pre-trained model, we introduce a multi-objective post-training task. We propose a composite reward that incorporates both intelligibility and prosody similarity, and adopt the Group Relative Policy Optimization (GRPO) algorithm [grpo] for training. Experimental results demonstrate that this post-training significantly enhances the model’s controllability with respect to text and melody, and also improves its generalization to out-of-pretraining-distribution data, such as instrumental sounds.
The contributions of this paper are summarized as follows:
  • We propose two audio tokenizers: (1) a unified music-notation-free prosody tokenizer capable of encoding prosody and melody from speech, singing voice, and even instrumental sounds; and (2) a unified timbre-disentangled content-style tokenizer that models linguistic content, prosody, and style for both speech and singing voice.
  • We introduce Vevo2, a unified and controllable framework for both speech and singing voice generation. Built upon the proposed speech-singing joint training, Vevo2 can be applied to versatile voice synthesis, conversion, and editing tasks by flexibly combining text, prosodic source, and style and timbre references during inference.
  • We propose a multi-objective post-training task for controllable voice generation. By optimizing both intelligibility and prosody similarity rewards with a GRPO-based algorithm, Vevo2 achieves greater stability and improved overall performance across a variety of tasks.
We will release code and model checkpoints of this study at Amphion [amphion]. [^2] A preliminary version of this work was released as a preprint on arXiv [vevo2-arxiv].
II Related Work
Controllable Voice Generation Controllable speech and singing voice generation has consistently remained a significant research focus. (1) In the speech domain, existing works explore control over multiple attributes: linguistic content, which is fundamental to TTS [tts-book-tanxu, xietianxin-survey] and speech editing tasks [voicebox, voicecraft, ssrspeech, f5tts]; timbre, which is central to VC tasks [ppg-vc, autovc, naturalspeech3, voiceshop, vevo, seedvc]; and style, which encompasses accent [parallel-ac-zhaoguanlong21, convertandspeak, vevo], emotion [esd, pavits, vevo], whisper [whisper-to-normal-lstm], and a broader natural language description [parler-tts, controlspeech, paraspeechcaps]. (2) For singing voice, beyond controlling lyrics, timbre, and style—as explored in SVS [diffsinger, visinger, tcsinger, stylesinger, techsinger], SVC [svcc-2023, svcc-vits-ziqian, neucosvc, amphion-svc], and singing lyric editing tasks [editsinger, unisinger, songeditor]melody control holds particular significance due to the inherent musicality of singing voices [musicgen, seedmusic, yue]. (3) From a unified modeling perspective, previous research has attempted to integrate various speech generation tasks such as TTS and VC [vevo, cosyvoice, metis], or different singing voice generation tasks like SVS and SVC [unisinger, everyone-can-sing]. However, models capable of unifying both speech and singing voice domains remain scarce [uniaudio, unisyn]. Moreover, the potential benefits of joint speech and singing pre-training, along with the versatile control mechanisms for multiple voice attributes within a unified system, remain largely unexplored.
Speech Tokenizer Recent research has emphasized the disentangled speech representation learning, aiming to separately model the factorized speech attributes to reduce learning complexity [megatts, naturalspeech3] and achieve more precise controllability [speech-resynthesis-interspeech21, naturalspeech3, vevo]. Common speech tokenizers include linguistic content tokenizers [hubert, vevo, syllablelm], content-style tokenizers [audiolm, seedtts, maskgct, cosyvoice, vevo], and prosody tokenizers [megatts, naturalspeech3]. However, in pursuing a unified controllable model for both speech and singing generation, existing speech tokenizers exhibit several limitations: As these tokenizers predominantly focus on the speech domain, their effectiveness in handling singing voice data remains unexplored. The design of a unified prosody tokenizer capable of handling both modalities, along with enhancing the melody encoding capabilities of content-style tokenizers [amphion-svc], remains an important issue for future exploration.
Post-training for Audio Generation Alignment via post-training has demonstrated its effectiveness in the generation of text [instructgpt, RLHF-anthoropic], vision [imagereward, tldr-perturb-vision], speech [speechalign, seedtts, cosyvoice2], music [musicrl], and sound effects [tango2, baton]. In speech generation, existing works have employed preference alignment to enhance multiple aspects of speech, including intelligibility [seedtts, cosyvoice2, intp], speaker similarity [seedtts, cosyvoice2], emotion controllability [seedtts, emo-dpo], and overall quality [speechalign, uno, fpo]. For developing a unified speech and singing voice generation model, the importance of specific capability elicitation through post-training becomes particularly crucial, given the expanded range of capabilities injected into the pre-trained model. However, the design of appropriate alignment objectives for unified speech and singing voice generation remains an open research problem.
(a) Prosody Tokenizer
(b) Content-Style Tokenizer
Figure 2: Unified speech and singing tokenizers of Vevo2.
III Methodology
The overall architecture of Vevo2 is a two-stage, versatile voice generation framework that unifies both speech and singing voice (Figure 1). During inference, by flexibly combining different input conditions—such as text, prosodic source, style reference, and timbre reference—Vevo2 supports a wide range of voice-related synthesis, conversion, and editing tasks (Section III-D). Specifically, Vevo2 first encodes linguistic, stylistic, and prosodic cues into discrete content-style tokens through an AR transformer, and then reconstructs Mel spectrograms via an FM transformer. These spectrograms are subsequently converted into waveforms using a Vocos-based vocoder [vocos]. To achieve the unified speech-singing voice generation and versatile controllability, Vevo2 is built upon several key methodological designs, which are detailed in the following sections.
Figure 3: Speech-Singing Joint Training with Explicit Prosody Learning (EPL) and Implicit Prosody Learning (IPL). We perform the next token prediction only on the sequence of content-style tokens (see Section III-B for more details).
III-A Unified Speech and Singing Voice Tokenizers
Prosody Tokenizer In order to model both the prosody of speech and the melody of singing voice using a unified method, we choose chromagram feature [meinard-book] and design a VQ-VAE tokenizer [vq-vae, rvqgan] based on it. Some previous studies of speech generation tried to use fundamental frequency (F0) [speech-resynthesis-interspeech21, naturalspeech3] or the low-frequency band of the mel spectrogram [megatts] to build a prosody tokenizer. The reasons why we adopt the chromagram are as follows:
  1. 1.
    Octave-free, thus friendly for unification: There is a significant distribution gap of F0 between speech and singing voice, such as the F0 range of singing voice is much wider [my-svc-paper-workshop]. Chromagram removes octave information between singing voice and speech, making it easier to capture common prosodic patterns across both domains.
  2. 2.
    Notation-free, thus friendly for scaling up: Chromagram effectively represents high-level musical and singing voice attributes such as melody and harmony [meinard-book], and can be automatically extracted from audio without manual transcription or music notation, making it suitable for large-scale data processing.
  3. 3.
    Robust for diverse audio domains: As a statistical distribution of spectral energy, chromagram is less sensitive to noise and pitch detection errors [yin], resulting in more stable performance. In addition, chromagram provides strong representations for both human voice and non-human sounds, such as instrumental music [meinard-book, musicgen]. These characteristics enhance the flexibility of melody control in downstream music generation tasks, such as enabling instrument-to-sing applications (Section V-F).